# LLMs’ Reshaping of People, Processes, Products, and Society in Software Development: A Comprehensive Exploration with Early Adopters

### v1

@@ 3.2.3. Implementation
  • Verifying through Manual Testing: Seven participants also engaged in manual testing practices, such as verifying code templates and solutions, playing around with isolated examples, and using developer tools or read–eval–print loop (REPL) tools to test code functionality before integration into their codebases. P10 shared that, for his process,“I verify its templates, and I verify its solutions. I put that in a inbox to test it.” Additionally, checking console logs, verifying expected function return values, and monitoring IDE warnings or errors were common practices among participants to ensure the correctness and integrity of ChatGPT-generated code. As P13 noted, “[I see] if my editor is like giving any warnings or errors.
  • Verifying through External Tests: Four participants employed external testing frameworks and tools, such as Pi Test (P9) and LangChain (P4), to evaluate the generated code. These tools facilitated automated testing, bug detection, and even collaboration between different agents to review and improve the code.
  • Verifying through External Sources: Two participants, one more experienced and one less, shared that they seek validation of their generated code quality from external sources such as the Internet. They look up documentation, search for common solutions on platforms like Stack Overflow, and review discussions to ensure that their implementation aligns with established practices and expectations. P13, a less-experienced developer, described it as an additional step she took for validating even after the code seemed to be functional: “I always take [the generated code] with a certain pinch of salt, and even though the code seems to work when I plug it in, I still always try and, you know, do a web search. […] I always go back and do a web search to ensure that, yeah, this is one of the common solutions to the problem, and it’s a good one. Basically reading all of the comments, like all of the discussions that happen on Stack Overflow, I think are really helpful to verify that my implementation is in line with how people would expect it to be done.
  • Clarifying Code through Generated Explanations: Two participants shared that they ask LLMs to explain their own generated code as part of their evaluation process; as P4 shared, he “[asked ChatGPT] to explain itself.” This involves requesting explanations for specific lines or segments of code in order to gain a deeper understanding of its logic, functionality, and rationale.
  • Assessing Visually and Logically: Two developers, who both worked in front-end development, noted that they visually assessed the quality of their code. P5 noted that he could both read and visually check the quality of his code: “I first run it in my head to see if it flows logically. […] So when I look at the code or look at the class, and I’d say, ‘Okay, that’s generally what I’m going for, based off of what I know.’ […] I’d either load up my own personal test environment and like a web browser, or I just copy and paste the code, and just see if it gives the result that I’m looking for.
  • Judging Quality through Readability: One participant, P7, shared how he used the process of assessing readability and adherence to established principles in order to judge the quality of the generated code. He shared that he “judge[d] whether it code is good or not” by evaluating the readability of the code through three strategies: applying the SOLID principles (Singh and Hassan, 2015); referring to the concepts he had learned through RC Martin’s clean-code manual Clean code: a handbook of agile software craftsmanship (Martin, 2009); and cross-referencing code against the programming language’s official documentation.
  • Improving via ”Self”-Correction: One participant, P12, shared his experience in which ChatGPT could recognize and rectify errors it had made in its code without explicit user guidance: “Sometimes [ChatGPT] literally will write something, or it’ll write some code, or even English text, and [you’ll] be like, ‘Hey? You made a mistake here. What was your mistake?’ And it will just correct it. You don’t even have to tell it what the mistake is. It’ll just be like, ‘Oh, my bad, I did this, this, this, and this wrong. Here, let me go fix it.’ And it’ll write it up for you.” This self-correction capability contributes to the reliability and usability of LLMs like ChatGPT in code generation tasks.
@@
  • Increased Skepticism: The existence of generated code from ChatGPT made five participants more skeptical and vigilant about evaluating the quality and reliability of the code. P1 shared that the existence of generated code made them more likely to check and test code: “[I]n a sense, yes, it has changed testing. I’m a bit more skeptical.
  • Necessity of Evaluation: ChatGPT has previously been found to have varying quality with regard to its generated code (Liu et al., 2023a, 2024). Seven participants emphasized the importance of thoroughly checking the generated code for accuracy and relevance. P4 mused on the general quality of generated code: “[Generated code is] not good enough that I don’t have to read it yet, although that might be nice one day— might put me out of a job, though.” Participants advised caution and suggested not solely relying on LLMs’ outputs without verification. P6 shared, “Be wary of what you get […] Don’t give up on checking the responses.” P1 particularly speculated on the hypothetical case in which ChatGPT’s generated code could pass tests but fail to fulfill the intended requirements. This underscores the necessity of thorough review processes, which should similarly be in-place for reviewing human-created code; as P16 shared, “[ChatGPT]’s not going to get it right every single time, and we need human overview on that […] And when it comes to human code. I think it’s the same thing.” Two participants shared that they paid more attention to reviewing the quality of generated code compared to human-written code, indicating a higher level of scrutiny for ChatGPT’s outputs.
  • Lack of Self-Evaluation One participant, P11, mentioned ChatGPT’s lack of self-evaluation as a limitation, which makes it inferior to even junior developers. He stated,“[…] It’s not evaluating its own code. It doesn’t actually have the ability to do those things that I would do to recode. I just can’t trust it in a certain way; I know it’s not making certain decisions based on facts or logic.” While this was one participant’s view, as noted in the above section on Evaluating Context, another participant (P12) commented that ChatGPT could correct itself.
  • Lack of Contextual Clarity: One participant, P4, emphasized the importance of context in understanding code, suggesting that humans have an advantage in having and being able to provide contextual information, especially about how and why code was created: “[I]f a human hands me a piece of code, I can ask them questions about it. If someone handed me code that was written by ChatGPT, I don’t know what I would want. […] would I want the prompt from the code? Because I’m missing the context now of how and why the code was generated.[…] If it’s a really small piece of code, I can probably just read it and figure it out and either choose to throw it away or go ask ChatGPT, ‘Hey, will you validate this code for me?’ I think if a human gave me ChatGPT code with the prompt that generated it, I would feel a lot better.
@@
  • Modifying Before Integration: Fifteen participants reported that they modified the generated code before integrating it into their projects. They described refining the code to align with their specific requirements and overcome potential errors. P13 shared her process: “Once I got that overall approach, I refined it in a few places, and then I sent it back to ChatGPT […] ’Kind of give me boilerplate code for that.’ So I think it did a really good job at giving me that boilerplate code. And then I just had to do a lot of refinements within it to overcome certain errors.
    P2 noted the need to adapt the logic of LLMs into the structure of existing projects and compared it to integrating code from other sources: ”[Integrating ChatGPT’s code results into my code base is] similar to what I used to do with Stack Overflow. It will definitely give me a snippet, but variable names and all of that will be very different. The code structure will be different. So I’ll just take the logic, but follow the structure which is already the pre-existing code base.”
  • Copy-Pasting Code: Twelve participants reported simply copy-pasting the generated code into their development environments. They found this process straightforward and efficient, allowing them to quickly incorporate the generated solutions into their projects. P7 stated, “Maybe there [aren’t] too [many] steps. Just asking the question, and if the answer looks good to me, I would just copy-paste.
  • Rewriting Code: Three participants mentioned that they preferred to rewrite the generated code manually in their development environments. They did this to gain a deeper understanding of the code and to mitigate potential errors that might arise from directly copy-pasting the generated solutions, giving them greater control over the code integration process. As P9 shared, “I re-implement [the code] in Python myself instead of just taking the output directly and pasting it there.
  • Using Generated Code: Participants exhibited varying degrees of frequency in discarding generated code. Six participants rarely discarded code, two disposed of it about half the time, and eight often discarded it. Those who seldom discarded code pointed out that, with well-crafted prompts and in certain programming languages, LLMs are capable of producing viable code, albeit with some limitations; as P13 shared, “I would say [I discard code] not very often. It seems to be useful in the context that I use it. I think I’ve been frequently throwing out code given by ChatGPT only in cases of SQL.
  • Discarding Generated Code: Eight participants shared that they threw away a lot (greater than 50%) of their generated code, while two shared that they threw away about half of their generated code. Frequent discards and reduced usage of generated code was more prevalent among those eight who used LLMs for ideation, consulting, or learning. P4 shared his reasons for throwing away code, along with the value that he still received from discarded code: “It’s good at getting me from zero to something, but I iterate a lot on it. I throw away a lot of code, and sometimes I choose not to do what it says to do, like maybe it’s using some design pattern, or it’s just being too clever with the code. I don’t generally have a problem with with how it’s doing stuff. It just might be different than how I would do it. For the type of stuff I’m building, I value being familiar with my own way more so than even code quality at times. […] I think it’s just the nature of the iterative prototyping that I do, that I throw away a lot of code, no matter what. But that doesn’t mean I don’t get value from the ChatGPT code. I think it still teaches me a lot of stuff.
@@
  • Friction with Copy-Pasting: One participant, P4, expressed his frustration with the copy-pasting process due to the need to switch between applications or rely on third-party plugins: “[E]very time you want to interact with ChatGPT, there is friction just because of the UI. You have to switch to the web application, or you have to get one of these unofficial third party plugins […] So right now I copy-paste, and I find that annoying, but the value is still high enough. So it’s like, what is the cost versus value gained.
3.2.4. Testing and Code Review
  • Generating Unit Tests: Seven participants had used LLMs for generating tests, particularly for unit tests due to their relatively basic and formulaic nature. As P9 shared, “Sometimes I use ChatGPT to create simple unit tests for me, instead of me writing from scratch again. But then, I only save those for smaller functions after I’ve done my refactoring.” One participant, P15, noted that using ChatGPT to generate tests encouraged him to incorporate more testing into his coding practices: “[ChatGPT] does write unit tests for me. So where I would not normally have them, if it’s able to write unit test for them, then I’ll have it do that. We don’t spend a lot of time creating tests as much as we should. But when when it comes to certain functions, like even just the basic functions […] usually [the generated tests] are pretty good when it comes to just a small function.
  • Simulating Code Reviews: Two participants, who worked independently as contractors or in small teams, used LLMs for code review due to the lack of colleagues to review their code; an additional participant used ChatGPT to review chunks of his code. For the two solo developers, using LLMs for code review served as a way to ensure their code met professional standards and to compensate for the absence of traditional code review processes within their teams. As P9 shared: “[R]ight now, I’m working as a contractor solo. I don’t have the privilege of getting somebody to do code reviews for me. So I would say in terms of code quality, that really helped me to maintain […] a professional coding level.
@@ 3.2.4. Testing and Code Review
  • Inapplicable for Code Review Eight participants indicated that they did not utilize ChatGPT for code reviews or that it had minimal impact on their code review process. P15, despite being a big proponent of using ChatGPT, expressed hesitance toward using it for code reviews: “I don’t use it for code review, mainly because I need to understand what the code is doing myself […] It’s better for me to know each step of what it’s trying to do, because I need to know … how it’s going to affect the rest of the system. So where it’s like, ‘Oh, we deleted this function chapter. [It]’s gonna say it’s fine, but in reality, that function’s being used in many places. […] That type of thing [the LLM] wouldn’t be able to do. It just doesn’t have the big picture.” P4 expressed significant concerns regarding using LLMs for code reviews and preferred human reviewers: “It would be silly to have [an LLM] replace a human, because one of the main benefits of doing a code review, at least in the teams that I work on, is transferring the knowledge amongst the developers. So using [an LLM] for [code reviews] would be a great way [sic] to remove the entire purpose of the code review.
  • Struggles with Complex Tests Two participants noted that LLMs struggled with generating larger or more system-wide tests. P2 noted that he did not consider using ChatGPT for system-wide testing due to security considerations. P15 shared further that ChatGPT lacked the context necessary to generate larger tests: “It gets a little bit harder if you were talking about […] an entire web page […] because it just doesn’t know what people are supposed to be doing [or] how it should create the test.
3.2.5. Debugging, Refactoring, and Documentation
  • Improving Debugging: LLM tools have shown promise in helping individuals debug code (Surameery and Shakor, 2023a; Haque and Li, 2023). Ten participants mentioned having used LLMs to debug their own code. Participants like P5 primarily utilized LLMs as an immediate debugging tool, relying on it to assist in troubleshooting specific issues with their code, especially in front-end development; as he shared, “I use [ChatGPT] as my immediate sort of debugging tool. So I would ask it particular questions [on a] front-end component that I’m trying to develop, and I use it to try to point me in the right direction.” For participants like P6, LLMs were especially useful in understanding bugs and their causes: “I always have questions about why something is failing, what something is doing. So [Copilot Chat] does help on a day-to-day basis with programs.
@@ 3.2.5. Debugging, Refactoring, and Documentation
    Participants highlighted LLMs’ efficacy in significantly reducing debugging time by providing quick solutions or guiding them in the right direction. P5 shared his experiences: “[ChatGPT] helps to dramatically shorten the whole debugging process. If it doesn’t give you the answer — that is, if it doesn’t give you the answer on the first try, […] it helps to put me in some right directions to where I can do some further research or ask it more questions.
  • Reducing Syntax-Based Errors: Two participants emphasized LLMs’ utility in resolving syntax-related errors, such as missing punctuation or braces. P14 stated that, “[W]henever I’m missing a comma somewhere or a brace. […] I just paste it into ChatGPT and say, ‘Fix the syntax mistake.’” He then shared how ChatGPT quickly identifies and fixes such errors, saving him from spending hours debugging simple syntax mistakes.
  • Performing Refactoring: We define refactoring as transforming the code in such a way as its functionality and behavior is preserved while improving its maintainability or comprehensibility (Fowler, 2018). Two participants shared that they used LLMs for refactoring purposes, aiming to enhance maintainability and comprehensibility while preserving functionality. P8 shared his process: “I have a method, a class, or something that I want to verify, to read, and I’m going to get an idea of how I can do that — refactor in a better way. […] [if] I agree with the refactoring that [ChatGPT] gave to me, I’ll paste it in my codebase. Otherwise, I try to only get the idea and implement it myself.” Additionally, one participant utilized LLMs to condense code, particularly by converting code blocks into more concise versions. P7 noted that he did so, sharing an example prompt and ChatGPT response: “’This is the code I [am] writing for a for loop. Can you just convert it to the stream version?’ And then [ChatGPT] just gave me this one line of code. So I feel this is pretty useful.
@@
  • Debugging Difficulties: One participant, P9, shared that he had had a negative experience while using ChatGPT to debug a piece of code he had found online, but with poor results. He emphasized that his lack of experience with the code left him with an inability to properly evaluate the results of ChatGPT’s output, leaving him frustrated:“I was using ChatGPT to help me debug and help me revise it. But since I don’t understand the code perfectly, and relying too much on ChatGPT at that point, it was giving me the incorrect prompt, which I didn’t know until two or three hours later […] I thought, ‘There’s something wrong with my input,’ and but it turns out it’s not. It’s actually ChatGPT — it was mixing up Python’s syntax in there. It was actually using syntax from all languages, but that almost looked like Python.
3.3. RQ3: How has the use of LLMs influenced the software products created?
Figure 4. Findings related to RQ3, emphasizing the pros and cons of LLMS with regard to their code-related output.
@@ 3.3. RQ3: How has the use of LLMs influenced the software products created?
3.3.1. Quality and Complexity of Generated Code
  • Producing Clean, Readable Code: In Section 3.1.2, we presented how some participants used LLMs to keep their code clean and neat, incorporate ideal design patterns, and understand industry best practices. Six participants shared that they found LLMs’ generated code to be clean, readable, and systematic. They likened its readability to that of standard Stack Overflow answers or official documentation for programming languages. Five participants, such as P4, appreciated the ease of understanding the syntax, which facilitated quick comprehension of the code: “[ChatGPT is] generally pretty good about not generating difficult to read code or overly complex [code]. And just asking it to improve the code in terms of readability generally works for small snippets. I’ve had good success with that.
  • Generating Code with Reasonable Complexity: Five participants stated that the code actually had fair to good complexity for the tasks that they were working on. They acknowledged that while LLMs may not always produce the most efficient solutions, the complexity generally aligned well with their needs.
    Particularly for common tasks or well-known problems, participants generally found the complexity to be sufficient. P11 stated that he found ChatGPT to have good complexity for more common tasks, but not for more novel tasks: “Every now and then I’ll put in a code challenge, or an interview kind of problem, and […] I think I think it does better on those, because other people have posted about those kinds of problems. And then you can ask it, ‘Oh, can you do this more efficiently? Can you do whatever?’ And it’ll talk to you and explain to you. But again, like I said, my suspicion is that it’s only because those are things that people talk about more. If I try to give it more unique data structure problems from my actual experience, it doesn’t do as well.
  • Conducting Complexity Analysis: Four participants utilized LLMs as tools for checking the time and space complexity of code snippets and explaining the benefits of different approaches. This served as a learning opportunity, particularly for participants who were newer to computing. P16 shared that she used ChatGPT to learn about code complexity, “[S]ometimes, if I don’t know, I’m just like [to ChatGPT], ‘Oh, what’s the time complexity, and the space complexity of this problem? And explain why and explain the benefits of using this method versus that method by its complexity. And [ChatGPT] does that sometimes. It’s like, ‘This saves space, but this saves time, and you can like solve this problem […]’
@@ 3.3.1. Quality and Complexity of Generated Code
  • Struggling with Up-to-Date Information: Two participants noted that ChatGPT’s generated code sometimes lacked up-to-date information, particularly when dealing with rapidly evolving technologies or APIs. This resulted in outdated or irrelevant suggestions, as P6 shared: “Sometimes it doesn’t work out well, in the sense that it just gives you suggestions that are a little outdated about the use of a specific API.
  • Over-Engineered, Complex Code: Five participants raised concerns about LLMs over-engineering solutions, adding unnecessary complexity and inefficiency to simple problems. They highlighted instances where LLMs introduced extraneous modules or convoluted solutions. For instance, as P8 complained, “One complaint that I have about ChatGPT’s output is over-engineering. […] In some of cases, I’m feeling that sometimes that I’m asking it, ‘Okay, write a simple method for, I don’t know, multiplying two numbers together.’ […] Sometimes it does the over-engineering for those cases.” P3 additionally shared this sentiment: “It can build some things, but you know, even that […] sometimes becomes more complicated than just coding it yourself, and you can’t create complex functionality.” For more novel or unique challenges, some participants (like P9 and P11) observed limitations in ChatGPT’s ability to produce code with optimal complexity.
3.3.2. Optimal Use Cases
  • Excelling at Small Tasks: Nine participants identified LLMs’ strength in handling small tasks, particularly those that involve routine or standard procedures like the boilerplate code referenced in Section 5. They found it most effective for tasks that could be decomposed to a low level, such as writing small code snippets or implementing basic functionalities. For ChatGPT, P4 shared that “I’ve kind of isolated its use cases to helping me improve code at the function level, like small snippets of code, and helping me go from zero to something”; this sentiment is connected to his relationship with frequently discarding code as part of an iterative process, as presented in Section 3.2.3.

### v2

@@ 4.2.3 Implementation
  • Verifying through Manual Testing: Seven participants also engaged in manual testing practices, such as verifying code templates and solutions, and using developer tools or REPL tools to test code functionality before integration. Checking console logs, verifying expected function return values, and monitoring IDE warnings or errors were common practices among participants.
  • Verifying through External Tests and Sources: Four participants employed external testing frameworks and tools to evaluate the generated code. Two participants sought validation from external sources such as the Internet, looking up documentation and searching platforms like Stack Overflow.
  • Other Verification Methods: Two participants asked LLMs to explain their own generated code as part of their evaluation process. Two developers working in front-end development noted that they visually assessed the quality of their code. One participant shared how he assessed readability using SOLID principles, clean code concepts, and official documentation. One participant noted ChatGPT’s self-correction capability in recognizing and rectifying errors without explicit user guidance.
@@
  • Increased Skepticism: Five participants became more skeptical and vigilant about evaluating code quality due to the existence of generated code. P1 shared: “[I]n a sense, yes, it has changed testing. I’m a bit more skeptical.
  • Necessity of Evaluation: Seven participants emphasized the importance of thoroughly checking the generated code for accuracy and relevance. P6 shared, “Be wary of what you get […] Don’t give up on checking the responses.” Two participants shared that they paid more attention to reviewing generated code compared to human-written code.
  • Lack of Self-Evaluation and Contextual Clarity: One participant mentioned ChatGPT’s lack of self-evaluation as a limitation. Another emphasized the importance of context in understanding code, suggesting that humans have an advantage in being able to provide contextual information about how and why code was created.
@@
  • Modifying Before Integration: Fifteen participants reported that they modified the generated code before integrating it into their projects. P13 shared her process: “Once I got that overall approach, I refined it in a few places, and then I sent it back to ChatGPT […] ’Kind of give me boilerplate code for that.’ So I think it did a really good job at giving me that boilerplate code. And then I just had to do a lot of refinements within it to overcome certain errors.
  • Copy-Pasting Code: Twelve participants reported simply copy-pasting the generated code into their development environments. P7 stated, “Maybe there [aren’t] too [many] steps. Just asking the question, and if the answer looks good to me, I would just copy-paste.
  • Rewriting Code: Three participants mentioned that they preferred to rewrite the generated code manually to gain a deeper understanding and mitigate potential errors.
  • Using and Discarding Generated Code: Six participants rarely discarded code, two discarded about half of all generated code, and eight discarded the majority of their generated code. Those who seldom discarded code pointed out that, with well-crafted prompts, LLMs are capable of producing viable code. Eight participants shared that they threw away a lot (greater than 50
@@
  • Friction with Copy-Pasting: One participant expressed frustration with the copy-pasting process due to the need to switch between applications or rely on third-party plugins.
4.2.4 Testing and Code Review
  • Generating Unit Tests: Seven participants had used LLMs for generating tests, particularly for unit tests due to their relatively basic and formulaic nature. One participant noted that using ChatGPT to generate tests encouraged him to incorporate more testing into his coding practices.
  • Simulating Code Reviews: Two participants, who worked independently as contractors or in small teams, used LLMs for code review due to the lack of colleagues to review their code; an additional participant used ChatGPT to review chunks of his code.
@@ 4.2.4 Testing and Code Review
  • Inapplicable for Code Review: Eight participants indicated that they did not utilize ChatGPT for code reviews or that it had minimal impact on their code review process. P15 expressed hesitance: “I don’t use it for code review, mainly because I need to understand what the code is doing myself […] It’s better for me to know each step of what it’s trying to do, because I need to know … how it’s going to affect the rest of the system.
  • Struggles with Complex Tests: Two participants noted that LLMs struggled with generating larger or more system-wide tests due to security considerations or lack of necessary context.
4.2.5 Debugging, Refactoring, and Documentation
@@
  • Improving Debugging: Ten participants mentioned having used LLMs to debug their own code. Participants highlighted LLMs’ efficacy in significantly reducing debugging time by providing quick solutions or guiding them in the right direction. P5 shared: “[ChatGPT] helps to dramatically shorten the whole debugging process. If it doesn’t give you the answer […] it helps to put me in some right directions to where I can do some further research or ask it more questions.
  • Reducing Syntax-Based Errors: Two participants emphasized LLMs’ utility in resolving syntax-related errors, such as missing punctuation or braces.
  • Performing Refactoring: Two participants shared that they used LLMs for refactoring purposes, aiming to enhance maintainability and comprehensibility while preserving functionality. Additionally, one participant utilized LLMs to condense code, particularly by converting code blocks into more concise versions.
@@
  • Debugging Difficulties: One participant shared a negative experience while using ChatGPT to debug code found online, emphasizing that his lack of experience with the code left him unable to properly evaluate ChatGPT’s output, leading to frustration when ChatGPT mixed up Python syntax with other languages.
4.3 RQ3: How has the use of LLMs influenced the software products created?
@@
4.3.1 Quality and Complexity of Generated Code
  • Producing Clean, Readable Code: Six participants shared that they found LLMs’ generated code to be clean, readable, and systematic. They likened its readability to that of standard Stack Overflow answers or official documentation for programming languages. Five participants, such as P4, appreciated the ease of understanding the syntax: “[ChatGPT is] generally pretty good about not generating difficult to read code or overly complex [code]. And just asking it to improve the code in terms of readability generally works for small snippets. I’ve had good success with that.
  • Generating Code with Reasonable Complexity: Five participants stated that the code actually had fair to good complexity for the tasks that they were working on. Particularly for common tasks or well-known problems, participants generally found the complexity to be sufficient. P11 stated that he found ChatGPT to have good complexity for more common tasks, but not for more novel tasks: “Every now and then I’ll put in a code challenge, or an interview kind of problem, and […] I think it does better on those, because other people have posted about those kinds of problems. […] If I try to give it more unique data structure problems from my actual experience, it doesn’t do as well.
  • Conducting Complexity Analysis: Four participants utilized LLMs as tools for checking the time and space complexity of code snippets and explaining the benefits of different approaches. This served as a learning opportunity, particularly for participants who were newer to computing.
@@ 4.3.1 Quality and Complexity of Generated Code
  • Struggling with Up-to-Date Information: Two participants noted that ChatGPT’s generated code sometimes lacked up-to-date information, particularly when dealing with rapidly evolving technologies or APIs.
  • Over-Engineered, Complex Code: Five participants raised concerns about LLMs over-engineering solutions, adding unnecessary complexity and inefficiency to simple problems. For instance, as P8 complained, “One complaint that I have about ChatGPT’s output is over-engineering. […] In some of cases, I’m feeling that sometimes that I’m asking it, ‘Okay, write a simple method for, I don’t know, multiplying two numbers together.’ […] Sometimes it does the over-engineering for those cases.” For more novel or unique challenges, some participants observed limitations in ChatGPT’s ability to produce code with optimal complexity.
4.3.2 Optimal Use Cases
  • Excelling at Small Tasks: Nine participants identified LLMs’ strength in handling small tasks, particularly those that involve routine or standard procedures. They found it most effective for tasks that could be decomposed to a low level, such as writing small code snippets or implementing basic functionalities. For ChatGPT, P4 shared that “I’ve kind of isolated its use cases to helping me improve code at the function level, like small snippets of code, and helping me go from zero to something”; this sentiment is connected to his relationship with frequently discarding code as part of an iterative process.